OpenStudy (filllip66):

The fourth power of an integer is x, and x is between 100 and 600. What is x

OpenStudy (phi):

I would test some numbers for example 2*2 * 2*2 i.e. 2^4 is 4*4 =16 which is too small try another guess

OpenStudy (solomonzelman):

(Don't want to give it away, sorry.)

OpenStudy (solomonzelman):

I will post an example.

OpenStudy (solomonzelman):

The 5th power of an integer is x, and x is between 30 and 1000. What is x? \(\color{black}{\displaystyle 2^5=32 }\) \(\color{black}{\displaystyle 4^5=1024 }\) So clearly the integer x is 3.

OpenStudy (filllip66):

so I just test numbers

OpenStudy (filllip66):

I got it it is 4

OpenStudy (phi):

what is 4*4*4*4 ?

OpenStudy (filllip66):

256

OpenStudy (phi):

yes, good job

OpenStudy (filllip66):

thanks

OpenStudy (filllip66):

oh yeah can you please tell me how to change your image

OpenStudy (phi):

go to your "detailed profile" and paste in a new picture

OpenStudy (filllip66):

where is detailed profile

OpenStudy (phi):

Hover/click over your name or icon.

OpenStudy (phi):

http://openstudy.com/users/filllip66